David's mass is 5kg less than john's, who in turn is 8kg lighter than than paul. If their total mass is 197kg, how heavy is each person?
Answer by TimothyLamb(4379)

x = david mass
y = john mass
z = paul mass
---
x = y - 5
y = z - 8
x + y + z = 197
---
put the system of linear equations into standard form
---
x - y = -5
y - z = -8
x + y + z = 197
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = david mass = 59.666667
y = john mass = 64.666667
z = paul mass = 72.666667
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php